Assuming there is no trade, if the Marginal Propensity to Consume is 0.80 and the government increases spending by $8 billion, by how much will output rise?

Select one:

a. By $24 billion

b. By $40 billion

c. By $48 billion

d. By $52 billion

e. By $60 billion
Answer #1

B

Change in output = Initial increase in expenditure * 1/MPS

Now MPS (Marginal propensity to save) = 1 -MPC = 1-0.8 = 0.2

Thus change in output = $8 billion/0.2 =   $40 billion